# Quantum algorithm for partial differential equations of non-conservative systems with spatially varying parameters

### v1

@@ III.1 Problem of interest
where $\dot{u}:=\partial u(t,\bm{x})/\partial t$. Note that we apply the operators in order from the right side; for instance, the product of $(2,1)$-component of the matrix and the first component of the vector on the right-hand side is calculated as $\sqrt{\kappa}\nabla((1/\sqrt{\varrho})\sqrt{\varrho}\dot{u})=\sqrt{\kappa}\nabla\dot{u}$. We omit dependence of variables on $(t,\bm{x})$ for simplicity. Although there are various mappings from the second-order PDE (13) to a first-order one, Eq. (16) has the following advantage. The key fact is that, as the dissipative nature of the target system, the norm of state vector of such equation must be reduced over time, and the reduction rate (will be shown to be $\|\bm{w}(0)\|/\|\bm{w}(T)\|$) directly represents the computational complexity for Hamiltonian simulation method; yet the reduction rate in the case of Eq. (16) is the minimum in a certain sense as proven in Appendix A. The PDE (14) can be simply represented as
$\frac{\partial u}{\partial t}$ $=-\frac{1}{\zeta}\left(-\nabla\cdot\kappa\nabla+\bm{\beta}\cdot\nabla+\alpha\right)u,\text{ for }(t,\bm{x})\in(0,T]\times\Omega.$ (17)
In the subsequent sections, we discuss how to spatially discretize Eqs. (16) and (17) followed by representing them as qubit systems.
@@ III.2.1 State vector and spatial difference operator
where $\sigma_{00}:=\ket{0}\!\bra{0}$ |, $\sigma_{11}:=\ket{1}\!\bra{1}$ |. Note again that for instance $\tilde{\varrho}$ represents a diagonal operator corresponding to $\varrho(\bm{x})$. Also, $B_{\mu}^{+}$ and $B_{\mu}^{-}$ denote the boundaries whose outward-pointing normal vectors are $\bm{e}^{[\mu]}$ and $-\bm{e}^{[\mu]}$, respectively. $\chi_{\mathrm{N}}$ and $\chi_{\mathrm{D}}$ are indicator functions defined as
@@
$\bm{A}$ $:=-\tilde{\zeta}^{-1}\left(\frac{1}{2}\sum_{\mu=0}^{d-1}\left(D_{\mu}^{+}\tilde{\kappa}D_{\mu}^{-}+D_{\mu}^{-}\tilde{\kappa}D_{\mu}^{+}\right)+\sum_{\mu=0}^{d-1}\left(\tilde{\beta}_{\mu}^{+}D_{\mu}^{-}+\tilde{\beta}_{\mu}^{-}D_{\mu}^{+}\right)+\tilde{\alpha}\right)$
$\quad+\frac{1}{2}\tilde{\zeta}^{-1}\sum_{\mu=0}^{d-1}\chi_{\mathrm{N}}(B_{\mu}^{-})D_{\mu}^{+}\tilde{\kappa}\left(I^{\otimes\sum_{\mu^{\prime}=\mu+1}^{d-1}}\otimes\sigma_{00}^{\otimes n_{\mu}}\otimes I^{\otimes\sum_{\mu^{\prime}=0}^{\mu-1}n_{\mu^{\prime}}}\right)$
$\quad-\frac{1}{2}\tilde{\zeta}^{-1}\sum_{\mu=0}^{d-1}\chi_{\mathrm{N}}(B_{\mu}^{+})D_{\mu}^{-}\tilde{\kappa}\left(I^{\otimes\sum_{\mu^{\prime}=\mu+1}^{d-1}}\otimes\sigma_{11}^{\otimes n_{\mu}}\otimes I^{\otimes\sum_{\mu^{\prime}=0}^{\mu-1}n_{\mu^{\prime}}}\right),$ (25)
where $\tilde{\beta}_{\mu}^{+}:=\max(\tilde{\beta}_{\mu},0)$ and $\tilde{\beta}_{\mu}^{-}:=\min(\tilde{\beta}_{\mu},0)$ for the upwind difference scheme [31], and the second and third lines of Eq. (III.2.1) appear due to the boundary condition. Note again that the discretization in Eq. (III.2.1) is just one example and we can use another finite difference scheme. The ODE $\mathrm{d}\bm{w}/\mathrm{d}t=-\bm{A}\bm{w}$ under the initial condition $w_{j}(0)=u(0,\bm{x}^{[j]})$ yields the solution of $w_{j}(t)=u(t,\bm{x}^{[j]})$. Since all qubit operators in the matrix $\bm{A}$ are represented by the tensor products of operators $\sigma_{01}$, $\sigma_{10}$, $\sigma_{00}$, $\sigma_{11}$, and $I$, we can calculate the multiplication of two operators in a symbolic manner based on some simple rules such as $\sigma_{01}\sigma_{10}=\sigma_{00}$, without any access to the full matrix. As we obtain the qubit operator representation of the matrix $\bm{A}$, in the next subsection we consider how to efficiently represent the diagonal operators in $\bm{A}$, such as $\tilde{\kappa}$ and $\tilde{\alpha}$.
III.2.2 Efficient quantum representation of spatially varying parameters
@@
We can reduce the number of terms by utilizing the relationship of $\sigma_{00}+\sigma_{11}=I$. Specifically, we aim to perform factorization such as $\sigma_{jj}\otimes\sigma_{00}+\sigma_{jj}\otimes\sigma_{11}=\sigma_{jj}\otimes(\sigma_{00}+\sigma_{11})=\sigma_{jj}\otimes I$. This example corresponds to the transformation of Boolean function $f=ab+a\bar{b}=a(b+\bar{b})=a$, where $a$ and $b$ are binary logic values. The problem of minimizing the number of terms in Eq. (III.2.2) corresponds to simplifying the Sum of Products (SOP) in Boolean algebra to represent certain logic function with fewer terms. A variety of methodologies have been developed to tackle this type of problem, such as Quine-McCluskey algorithm (QMC) [32, 33] and Espresso algorithm (EA) [34, 35]. QMC searches all candidate of the solution, hence it has the exact solution capabilities albeit with exponential computational complexity [36]. Although some improved methods in scaling are developed [37], it is suitable for small-scale problems. On the other hand, EA efficiently search for candidate solutions by eliminating redundancy, hence it is particularly efficient for scaling. Although the analytical evaluation of the compression rates, the number of variables, and computational complexity are not clarified, the empirical method have demonstrated that highly accurate approximate solutions can be obtained with small computational cost, as evidenced in various practical results [38]. In this study, we employed EA, considering the balance between execution time and optimality. Here, note the fact that our problem setting is not written by Boolean algebra, but each terms have continuous valued coefficients. As far as Boolean algebra, the overlap of each terms is not an issue since the final output value is calculated by summing up all terms with the logical OR operation. However, in our current problem setting, the overlap of different terms could cause the duplications of coefficient for a certain coordinate. To avoid this, we have added a process to check for duplicate coordinates in the compressed logic obtained using EA. The specific workflow is described in Algorithm 1. Note that this process for resolving duplicate coordinates can also be implemented efficiently, i.e., $O(|\mathcal{I}|^{2}n)$. It is also worth noting that the aforementioned procedure can be easily extended to the case where a spatially varying parameter $c(\bm{x})$ is a piece-wise constant function.
@@
III.3 Coefficient oracle by matrix product state
Here, we discuss how to implement the LCU coefficient oracle $O_{\mathrm{coef}}$ in efficient and approximate manners via matrix product states (MPSs) [39, 40]. Suppose that the integer $a$ is written in $n_{\mathrm{a}}$-bits binary representation as $a=(a_{n_{\mathrm{anc}}-1}\dots a_{0})_{2}$, where $a_{m}\in\{0,1\}$ is the $m$-th bit of the binary representation of $a$. Here, we discretize the integration range of $k$ in Eq. (3) into $2^{n_{\mathrm{a}}}$ integration points $\{k_{a}\}_{a=0}^{2^{n_{\mathrm{a}}}-1}$ ranging from $-2^{n_{\mathrm{a}}-n_{\mathrm{frac}}-1}$ to $22^{n_{\mathrm{a}}-n_{\mathrm{frac}}}-$ 2, at intervals of $2^{-n_{\mathrm{frac}}}$ where $n_{\mathrm{frac}}\in\mathbb{N}$. To this end, we define the integration point $k_{a}$, as follows:
$k_{a=(a_{n_{\mathrm{anc}}-1}\dots a_{0})_{2}}:=\left(-a_{n_{\mathrm{a}}-1}2^{n_{\mathrm{a}}-1}+\sum_{m=0}^{n_{\mathrm{a}}-2}a_{m}2^{m}\right)2^{-n_{\mathrm{frac}}}.$ (27)
This representation corresponds to associating $k_{a}$ with the binary $(a_{n_{\mathrm{a}}-1}\dots a_{0})_{2}$ regarded as the signed fractional binary with the integer part of $n_{\rm a}-n_{\rm frac}-1$ bits and the fractional part of $n_{\rm frac}$ bits. If an MPS corresponding to a quantum state $(1/\sqrt{\|\bm{c}\|_{1}})\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sqrt{c_{a}}\ket{a}$ is obtained, a quantum circuit for preparing this quantum state can be constructed [41, 42]. Here, we set $\omega_{a}=2^{-n_{\mathrm{frac}}}$, i.e., $c_{a}=2^{-n_{\mathrm{frac}}}/\pi(1+k_{a}^{2})$ in the first step described in Section II.1, which corresponds to the rectangular approximation of the integral with respect to $k$ in Eq. (3). To prepare $(1/\sqrt{\|\bm{c}\|_{1}})\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sqrt{c_{a}}\ket{a}$ in the form of an MPS, we first treat $[k_{0}\dots k_{2^{n_{\mathrm{a}}}-1}]^{\top}$ as a $2^{n_{\mathrm{a}}}$-dimensional vector and represent it as an MPS as follows:
$\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}k_{a}\ket{a}=\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sum_{b_{n_{\mathrm{a}}}=1}^{n_{\mathrm{a}}}\dots\sum_{b_{1}=1}^{n_{\mathrm{a}}}K_{a_{n_{\mathrm{a}}}b_{n_{\mathrm{a}}}}^{[n_{\mathrm{a}}]}K_{b_{n_{\mathrm{a}}}a_{n_{\mathrm{a}}-1}b_{n_{\mathrm{a}}-1}}^{[n_{\mathrm{a}}-1]}\dots K_{b_{3}a_{2}b_{2}}^{[2]}K_{b_{2}a_{1}}^{[1]}\ket{a_{n_{\mathrm{a}}}a_{n_{\mathrm{a}}-1}\dots a_{1}},$ (28)

### v2

@@ III.1 Problem of interest
where $\dot{u}:=\partial u(t,\bm{x})/\partial t$. Note that we apply the operators in order from the right side; for instance, the product of $(2,1)$-component of the matrix and the first component of the vector on the right-hand side is calculated as $\sqrt{\kappa}\nabla((1/\sqrt{\varrho})\sqrt{\varrho}\dot{u})=\sqrt{\kappa}\nabla\dot{u}$. We omit dependence of variables on $(t,\bm{x})$ for simplicity. Although there are various mappings from the second-order PDE (15) to a first-order one, Eq. (18) has the following advantage. The key fact is that, as the dissipative nature of the target system, the norm of state vector of such equation must be reduced over time, and the reduction rate (will be shown to be $\|\bm{w}(0)\|/\|\bm{w}(T)\|$) directly represents the computational complexity for Hamiltonian simulation method; yet the reduction rate in the case of Eq. (18) is the minimum in a certain sense as proven in Appendix A. Specifically, the norm reduction is governed only by the damping factor $\zeta(\bm{x})$. For example, the acoustic wave equation we will focus on later in Eq. (43) falls into Eq. (15) by setting $\zeta(\bm{x})=0$, $\kappa(\bm{x})=1$, $\alpha(\bm{x})=0$, and $\rho(\bm{x})=1/c(\bm{x})^{2}$ with $c(\bm{x})$ the spatially varying speed of sound. At that moment, the boundary condition $\bm{n}(\bm{x})\cdot\nabla u(t,\bm{x})=0$ corresponds to the sound hard boundary condition, i.e., the wall boundary, while $u(t,\bm{x})=0$ corresponds to the sound soft boundary condition, i.e., the free end. Since the acoustic equation has no damping factor $(\zeta(\bm{x})=0)$, it can be encoded in Eq. (18) without any norm reduction. The PDE (16) can be simply represented as
$\frac{\partial u}{\partial t}$ $=\left(\nabla\cdot\kappa\nabla-\bm{\beta}\cdot\nabla-\alpha\right)u,\text{ for }(t,\bm{x})\in(0,T]\times\Omega.$ (19)
The reduction rate of the norm, which directly affects the computational complexity, is evaluated in Appendix B. Specifically, the norm reduction is governed by the diffusion and absorption coefficients $\kappa(\bm{x})$ and $\alpha(\bm{x})$ under the conservation condition of $\bm{\beta}(\bm{x})$. For example, the heat equation we will focus on later in Eq. (53) falls into Eq. (16) by setting $\beta(\bm{x})=0$ and $\alpha(\bm{x})=0$. At that moment, the boundary condition $\bm{n}(\bm{x})\cdot\kappa(\bm{x})\nabla u(t,\bm{x})=0$ corresponds to the thermal insulation, while $u(t,\bm{x})=0$ corresponds to the prescribed temperature. Since the heat equation has the diffusion coefficient $(\kappa(\bm{x})>0)$, it results in the norm reduction depending on the magnitude of the coefficient. In the subsequent sections, we discuss how to spatially discretize Eqs. (18) and (19) followed by representing them as qubit systems.
@@ III.2.1 State vector and spatial difference operator
where $\sigma_{00}:=\ket{0}\!\bra{0}$ |, $\sigma_{11}:=\ket{1}\!\bra{1}$ |. Note again that for instance $\tilde{\varrho}$ represents a diagonal operator corresponding to $\varrho(\bm{x})$. Also, $B_{\mu}^{+}$ and $B_{\mu}^{-}$ denote the boundaries whose outward-pointing normal vectors are $\bm{e}_{\mu}$ and $-\bm{e}_{\mu}$, respectively. $\chi_{\mathrm{N}}$ and $\chi_{\mathrm{D}}$ are indicator functions defined as
@@
$\bm{A}$ $:=-\left(\frac{1}{2}\sum_{\mu=0}^{d-1}\left(D_{\mu}^{+}\tilde{\kappa}D_{\mu}^{-}+D_{\mu}^{-}\tilde{\kappa}D_{\mu}^{+}\right)+\sum_{\mu=0}^{d-1}\left(\tilde{\beta}_{\mu}^{+}D_{\mu}^{-}+\tilde{\beta}_{\mu}^{-}D_{\mu}^{+}\right)+\tilde{\alpha}\right)$
$\quad+\frac{1}{2}\sum_{\mu=0}^{d-1}\chi_{\mathrm{N}}(B_{\mu}^{-})D_{\mu}^{+}\tilde{\kappa}\left(I^{\otimes\sum_{\mu^{\prime}=\mu+1}^{d-1}}\otimes\sigma_{00}^{\otimes n_{\mu}}\otimes I^{\otimes\sum_{\mu^{\prime}=0}^{\mu-1}n_{\mu^{\prime}}}\right)$
$\quad-\frac{1}{2}\sum_{\mu=0}^{d-1}\chi_{\mathrm{N}}(B_{\mu}^{+})D_{\mu}^{-}\tilde{\kappa}\left(I^{\otimes\sum_{\mu^{\prime}=\mu+1}^{d-1}}\otimes\sigma_{11}^{\otimes n_{\mu}}\otimes I^{\otimes\sum_{\mu^{\prime}=0}^{\mu-1}n_{\mu^{\prime}}}\right)$
$\quad+\frac{1}{2}\sum_{\mu=0}^{d-1}\chi_{\mathrm{D}}(B_{\mu}^{-})\tilde{\kappa}\left(I^{\otimes\sum_{\mu^{\prime}=\mu+1}^{d-1}}\otimes\sigma_{00}^{\otimes n_{\mu}}\otimes I^{\otimes\sum_{\mu^{\prime}=0}^{\mu-1}n_{\mu^{\prime}}}\right)$
$\quad+\frac{1}{2}\sum_{\mu=0}^{d-1}\chi_{\mathrm{D}}(B_{\mu}^{+})\tilde{\kappa}\left(I^{\otimes\sum_{\mu^{\prime}=\mu+1}^{d-1}}\otimes\sigma_{11}^{\otimes n_{\mu}}\otimes I^{\otimes\sum_{\mu^{\prime}=0}^{\mu-1}n_{\mu^{\prime}}}\right),$ (27)
where $\tilde{\beta}_{\mu}^{+}:=\max(\tilde{\beta}_{\mu},0)$ and $\tilde{\beta}_{\mu}^{-}:=\min(\tilde{\beta}_{\mu},0)$ for the upwind difference scheme [34], and the second and subsequent lines of Eq. (III.2.1) appear due to the boundary condition. Note again that the discretization in Eq. (III.2.1) is just one example and we can use another finite difference scheme. The ODE $\mathrm{d}\bm{w}/\mathrm{d}t=-\bm{A}\bm{w}$ under the initial condition $w_{j}(0)=u(0,\bm{x}^{[j]})$ yields the solution of $w_{j}(t)=u(t,\bm{x}^{[j]})$. Since all qubit operators in the matrix $\bm{A}$ are represented by the tensor products of operators $\sigma_{01}$, $\sigma_{10}$, $\sigma_{00}$, $\sigma_{11}$, and $I$, we can calculate the multiplication of two operators in a symbolic manner based on some simple rules such as $\sigma_{01}\sigma_{10}=\sigma_{00}$, without any access to the full matrix. As we obtain the qubit operator representation of the matrix $\bm{A}$, in the next subsection we consider how to efficiently represent the diagonal operators in $\bm{A}$, such as $\tilde{\kappa}$ and $\tilde{\alpha}$. As we will see later, we actually need an assumption that the spatially varying parameters are piecewise constant functions so that the resulting qubit operators could be represented efficiently.
III.2.2 Efficient operator representation of spatially varying parameters
@@
We can reduce the number of terms by utilizing the relationship of $\sigma_{00}+\sigma_{11}=I$. Specifically, we aim to perform factorization such as $\sigma_{jj}\otimes\sigma_{00}+\sigma_{jj}\otimes\sigma_{11}=\sigma_{jj}\otimes(\sigma_{00}+\sigma_{11})=\sigma_{jj}\otimes I$. This example corresponds to the transformation of Boolean function $f=ab+a\bar{b}=a(b+\bar{b})=a$, where $a$ and $b$ are binary logic values. The problem of minimizing the number of terms in Eq. (III.2.2) corresponds to simplifying the Sum of Products (SOP) in Boolean algebra to represent a certain logic function with fewer terms.
@@
A variety of methodologies have been developed to tackle this type of problem, such as Quine-McCluskey algorithm (QMC) [35, 36] and Espresso algorithm (EA) [37, 38]. QMC searches all candidate of the solution, hence it has the exact solution capabilities albeit with exponential computational complexity [39]. Although some improved methods in scaling are developed [40], it is suitable for small-scale problems. On the other hand, EA efficiently search for candidate solutions by eliminating redundancy, hence it is particularly efficient for scaling. Although the analytical evaluation of the compression rates, the number of variables, and computational complexity are not clarified, the empirical method have demonstrated that highly accurate approximate solutions can be obtained with small computational cost, as evidenced in various practical results [41]. In this study, we employed EA, considering the balance between execution time and optimality.
Figure 2: Concept of the simplification of operators for spatially varying parameters. (a) Example of spatially varying parameters on a $4\times 4$ grid where black pixels have $c_{1}=1$ and white pixels have $c_{0}=0$. $x$ and $y$ coordinates corresponds to the bit strings $b_{1}b_{0}$ and $b_{3}b_{2}$, respectively. (b) Truth table of the Boolean function $f(b_{3},b_{2},b_{1},b_{0})$ which takes the value of $1$ when the corresponding pixel is black. (c) Diagonal operator corresponding to the spatially varying parameters given in (a) before simplification. (d) Truth table of the Boolean function $f(b_{3},b_{2},b_{1},b_{0})$ after simplifying (b) by the logic minimization technique. (e) Diagonal operator corresponding to the spatially varying parameters given in (a) after simplification.
Figure 2 illustrates the conceptual diagram of simplification of the operators for spatially varying parameters by a logic minimization technique. Let us consider the spatially varying parameters on a $4\times 4$ grid where black pixels have the value of $c_{1}=1$ while white pixels have the value of $c_{0}=0$ as shown in Fig. 2(a). Associating the coordinate and the value of $c_{1}$ and $c_{0}$ with bit strings and binary values, respectively, we can write a truth table as shown in Fig. 2(b). This truth table corresponds to the diagonal operator in Fig. 2(c), which is exactly Eq. (III.2.2). Applying the logic minimization technique, specifically EA, for the Boolean function $f(b_{3},b_{2},b_{1},b_{0})$, we obtain the simplified Boolean function as shown in Fig. 2(d) where the symbol -- means that the output does not depend on the input variable, i.e., $b_{2}$. This simplified truth table corresponds to the diagonal operator in Fig. 2(e), which is simplified from Fig. 2(c). In this way, we can obtain the efficient (simplified) diagonal operator representation of spatially varying parameters by applying the EA for the truth table corresponding naive representation of spatially varying parameters.
Here, note the fact that our problem setting is not written by Boolean algebra, but each terms have continuous valued coefficients. As far as Boolean algebra, the overlap of each terms is not an issue since the final output value is calculated by summing up all terms with the logical OR operation. However, in our current problem setting, the overlap of different terms could cause the duplications of coefficient for a certain coordinate. To avoid this, we have added a process to check for duplicate coordinates in the compressed logic obtained using EA. The specific workflow is described in Algorithm 1. Note that this process for resolving duplicate coordinates can also be implemented efficiently, i.e., $O(|\mathcal{I}|^{2}n)$. It is also worth noting that the aforementioned procedure can be easily extended to the case where a spatially varying parameter $c(\bm{x})$ is a piecewise constant function. That is, if the spatially varying parameters are piecewise constant functions, the logic minimizer technique is applicable for efficient representation of the operators. Therefore, we assume in the present study that the spatially varying parameters are piecewise constant functions. This assumption is reasonable in the engineering application because spatially varying parameters usually reflect the material properties in the computational domain, which makes spatially varying parameters be piecewise constant. For instance, if we have two materials in the computational domain, spatially varying parameters take two values each of which corresponds to each material property.
@@
III.3 Coefficient oracle realized using matrix product state
Here, we discuss how to implement the LCU coefficient oracle $O_{\mathrm{coef}}$ in efficient and approximate manners via matrix product states (MPSs) [42, 43]. Suppose that the integer $a$ is written in $n_{\mathrm{anc}}$-bits binary representation as $a=(a_{n_{\mathrm{anc}}-1}\dots a_{0})_{2}$, where $a_{m}\in\{0,1\}$ is the $m$-th bit of the binary representation of $a$. Here, we discretize the integration range of $k$ in Eq. (3) into $2^{n_{\mathrm{anc}}}$ integration points $\{k_{a}\}_{a=0}^{2^{n_{\mathrm{anc}}}-1}$ ranging from $-2^{n_{\mathrm{anc}}-n_{\mathrm{frac}}-1}$ to $22^{n_{\mathrm{anc}}-n_{\mathrm{frac}}}-$ 2, at intervals of $2^{-n_{\mathrm{frac}}}$ where $n_{\mathrm{frac}}\in\mathbb{N}$. To this end, we define the integration point $k_{a}$, as follows:
$k_{a=(a_{n_{\mathrm{anc}}-1}\dots a_{0})_{2}}:=\left(-a_{n_{\mathrm{anc}}-1}2^{n_{\mathrm{anc}}-1}+\sum_{m=0}^{n_{\mathrm{anc}}-2}a_{m}2^{m}\right)2^{-n_{\mathrm{frac}}}.$ (29)
This representation corresponds to associating $k_{a}$ with the binary $(a_{n_{\mathrm{anc}}-1}\dots a_{0})_{2}$ regarded as the signed fractional binary with the integer part of $n_{\rm a}-n_{\rm frac}-1$ bits and the fractional part of $n_{\rm frac}$ bits. If an MPS corresponding to a quantum state $(1/\sqrt{\|\bm{c}\|_{1}})\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sqrt{c_{a}}\ket{a}$ is obtained, a quantum circuit for preparing this quantum state can be constructed [44, 45, 46, 47]. Here, we set $\omega_{a}=2^{-n_{\mathrm{frac}}}$, i.e., $c_{a}=2^{-n_{\mathrm{frac}}}/\pi(1+k_{a}^{2})$ in the first step described in Section II.1, which corresponds to the rectangular approximation of the integral with respect to $k$ in Eq. (3). To prepare $(1/\sqrt{\|\bm{c}\|_{1}})\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sqrt{c_{a}}\ket{a}$ in the form of an MPS, we first treat $[k_{0}\dots k_{2^{n_{\mathrm{anc}}}-1}]^{\top}$ as a $2^{n_{\mathrm{anc}}}$-dimensional vector and represent it as an MPS as follows:
$\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}k_{a}\ket{a}=\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sum_{b_{n_{\mathrm{anc}}}=1}^{n_{\mathrm{anc}}}\dots\sum_{b_{1}=1}^{n_{\mathrm{anc}}}K_{a_{n_{\mathrm{anc}}}b_{n_{\mathrm{anc}}}}^{[n_{\mathrm{anc}}]}K_{b_{n_{\mathrm{anc}}}a_{n_{\mathrm{anc}}-1}b_{n_{\mathrm{anc}}-1}}^{[n_{\mathrm{anc}}-1]}\dots K_{b_{3}a_{2}b_{2}}^{[2]}K_{b_{2}a_{1}}^{[1]}\ket{a_{n_{\mathrm{anc}}}a_{n_{\mathrm{anc}}-1}\dots a_{1}},$ (30)